OpenStudy (anonymous):

Solve and show work for . . . . 1/2n-17/12=2/3n-13/6

OpenStudy (anonymous):

n=9/2

OpenStudy (anonymous):

how

OpenStudy (anonymous):

1/2n-17/12=2/3n-13/6 First find a common denominator which will be 12 6/12n-17/12=8/12n-26/12 add 26/12 to both sides 6/12n+9/12=8/12n subtract 6/12n from both sides 9/12=2/12n divide both sides by 2/12 n=9/2

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

good deal!